# Programmable Cognitive Bias in Social Agents

### v2

@@ Abstract
This paper introduces CoBRA, a novel toolkit for systematically specifying agent behavior in LLM-based social simulation. We found that conventional approaches that specify agent behaviors through implicit natural language descriptions cannot yield consistent behaviors across models, and the produced agent behaviors do not capture the nuances of the descriptions. In contrast, CoBRA presents a new approach to program agents’ cognitive biases explicitly, by grounding agents’ expected behaviors using classic social science experiments. CoBRA has two components: (1) Cognitive Bias Index that measures the cognitive bias of a social agent, by quantifying the agent’s reactions in a set of validated classical social science experiments; (2) Behavioral Regulation Engine that aligns the agent’s behavior to demonstrate controlled cognitive bias. We evaluated CoBRA as an HCI toolkit through demonstration and technical benchmarks. Our results suggest that CoBRA can precisely program the cognitive bias demonstrated in a social agent in a model-agnostic manner.
@@
Figure 1: Existing social simulation experiments often use implicit natural language descriptions to specify agent behaviors. However, we found that these specifications often resulted in inconsistent and unpredictable agent behaviors. For example, Ⓐ real-world economists are supposed to be less susceptible to the Framing Effect than the general population (Thaler & Sunstein, 2008); however, Ⓑ agents based on implicit natural language specifications often produce inconsistent behaviors across models, and the expected differences in behavior across roles (e.g., economists being less prone than laypeople) are not reliably observed. To tackle this challenge, Ⓒ we introduce CoBRA  which enables researchers to explicitly specify the cognitive biases of LLM-based agents quantitatively, thereby producing precise and consistent behaviors across models.
1 Introduction
@@
Figure 2: Example closed-loop workflow of CoBRA. A social scientist aims to create an agent with a moderate framing effect (e.g., 2.6 on a 0–4 scale). ① She specifies the desired bias level in CoBRA alongside the natural language agent description. ② CoBRA measures the agent’s framing effect using validated classical social science experiments (e.g., the Asian Disease study (Tversky & Kahneman, 1981)). ③ If the measured bias deviates from the specification, the Behavioral Regulation Engine iteratively adjusts the agent—through prompt engineering, activation modifications, or fine-tuning—until the agent reliably demonstrates the target bias.
@@
The simulation of complex human behaviors through LLM-based agents is rapidly emerging as a powerful tool for social scientists (Bail, 2024; Park et al., 2023), offering new ways to test social theories (Anthis et al., 2025a; Wu et al., 2025) and explore sensitive or hard-to-study behaviors without real-world consequences (Hwang et al., 2025). A critical component of these simulations is specifying agent behaviors (Yang et al., 2025; Chandramouli et al., 2024; Goyal et al., 2024) to control how the agent responds to prompts. For example, Park et al. specified the attributes of one of the agents in their paper using the following natural language description: ”John Lin is a pharmacy shopkeeper at … who loves to help people. He is always looking for ways to make the process of getting medication easier for his customers…” (Park et al., 2023).
@@
These implicit natural language descriptions have become the dominant approach for specifying social agent behaviors and are used in nearly all social simulation experiments (Rath et al., 2025; Park et al., 2023; Zhou et al., 2024; Li et al., 2023a; Zheng et al., 2023; Chen et al., 2024b; Jia et al., 2024). In the descriptions, researchers often specify demographic descriptors (e.g., age, occupation, family structure), personality traits (e.g., Big Five (Gosling et al., 2003) or MBTI (Furnham, 1996)), personal interests (e.g., hobbies, preferences), social relationships (e.g., friends, family, colleagues), and professional or task-oriented identities (e.g., judge, programmer). These practices rest on a growing understanding that LLMs, having been trained on vast text corpora, possess both perceptual skills (Wang et al., 2024b; Liu et al., 2024) that enable them to detect subtle cues in implicit language (Gallegos et al., 2024), and strong role-playing capabilities (Park et al., 2023)—drawing on patterns that link behaviors to human characteristics (Fast et al., 2016; Shao et al., 2023b).
@@
However, the research community has little empirical understanding of whether this approach is sufficient for reliable social simulations. For example, would the conclusion of the social simulation change if we change the foundation models, the prompts, or the temperatures? As a consequence, the conclusions of most LLM-based social simulations often are broad qualitative impressions—such as noting that a simulation is plausible and exhibits an echo chamber similar to the real world (Wang et al., 2024a; Zheng & Tang, 2024).
@@
In this paper, we first conducted a pilot experiment to empirically examine the limitations of implicit natural language specification (Section 3). We designed four agent specifications by adapting examples from prior work (Chen et al., 2024b; Park et al., 2023), implemented them across four foundation models, and evaluated their responses using a classic social science experiment, the Asian Disease study (Tversky & Kahneman, 1981). The results show that (1) the same specification produced inconsistent behaviors across models, and (2) implicit specification did not reliably produce the expected behaviors (e.g., trained experts like economists are less prone to the framing effect (List, 2003)).
@@
We then developed CoBRA—the Cognitive Bias Regulator for Social Agents (Figure 1)—a novel toolkit that enables users to precisely program the cognitive biases exhibited by social agents across different models. Compared to implicit natural language specifications, CoBRA offers two key advantages. First, it enhances reproducibility of agent behaviors across models, which is critical given that access to specific foundation models may change over time (Chen et al., 2024a). Second, it provides explicit and quantitative control over agent behaviors, rather than relying on unpredictable model behaviors (Sclar et al., 2023). For example, a scientist can directly specify that an agent should be less susceptible to framing effects.
@@
At the heart of CoBRA is a novel closed-loop system primitive that continuously measures and regulates agents’ cognitive biases by grounding their behaviors in classic social science experiments. Classic experiments provide validated and structured protocols: they define precise stimuli, expected behavioral distributions, and well-documented bias effects. For instance, the Asian Disease study (Tversky & Kahneman, 1981) has been extensively replicated and is widely regarded as a canonical demonstration of the framing effect: when identical outcomes are framed in terms of lives saved versus lives lost, human participants systematically shift between risk-averse and risk-seeking choices. Within CoBRA, such experiments function as calibration tasks: the system first measures whether an agent demonstrates the expected bias by observing its responses in these experiments and then adjusts the agent’s specification to align it with the target behavioral profile (Figure 2). We further hypothesize that aligning an agent’s behavior with target profiles in a set of classical experiments will reliably induce the corresponding cognitive biases, which will then generalize and manifest across a broader range of tasks (Section 7.2).
@@
Our implementation of CoBRA has two components: (1) Cognitive Bias Index that measures the cognitive bias of a social agent by quantifying the agent’s reactions in a set of validated classical social science experiments; (2) Behavioral Regulation Engine that aligns the agent’s behavior to demonstrate controlled cognitive bias. The Cognitive Bias Index contains eight classic social experiments mapped to four representative biases—Authority Effect, Bandwagon Effect, Confirmation Bias, and Framing Effect. The Behavioral Regulation Engine employs three approaches (prompt engineering (Rath et al., 2025; Zhuo et al., 2024), representation engineering (Zou et al., 2023; Rimsky et al., 2024; Todd et al., 2024), and fine-tuning (Ouyang et al., 2022; Rafailov et al., 2023; Hu et al., 2022; Ilharco et al., 2023)) to align agent behaviors.
@@
We position CoBRA as an HCI toolkit (Nebeling, 2017) that empowers social scientists to create new types of social simulations. We evaluated CoBRA using two common strategies for assessing HCI toolkits (Ledo et al., 2018): Technical Benchmark and Demonstration. In the benchmark, we show that (1) the cognitive bias induced through one classical experiment generalizes to other tasks (Section 7.2); (2) agents specified through CoBRA exhibit consistent behavioral tendencies regardless of the foundation model, sampling temperature, or reasoning mode (Section 7.3); and (3) CoBRA enables precise control under both open-weight and API-only settings (Section 7.4). In the Demonstration, we show how social scientists can use CoBRA to simulate the Emotional Contagion experiment (Kramer et al., 2014), where the participant’s social influence susceptibility (i.e., bandwagon effect bias) can be manipulated (Section 8). The simulated experiment validates a classical social science theory: agents exposed to varying levels of negative posts exhibited predictable sentiment shifts based on their bias levels programmed by CoBRA, with higher Bandwagon Effect levels leading to stronger emotional contagion.
@@
This work makes four main contributions. First, we present the first empirical evidence that implicit natural language specifications, the dominant method for defining LLM-based social agents, fail to produce reliable and reproducible behaviors across models. Second, we introduce CoBRA, a novel HCI toolkit that enables social scientists to precisely program and control cognitive biases in agents, improving reproducibility and allowing explicit, quantitative alignment. Third, we contribute a closed-loop system primitive that grounds agent behaviors in validated and structured social science experiments, using them as primitives for both measuring and regulating cognitive biases. Finally, we demonstrate CoBRA through technical benchmarks and a case study of the Emotional Contagion experiment, showing that induced biases generalize across tasks, remain robust across models and settings, and allow social scientists to create new forms of controlled social simulations.
2 Background & Related Work
This section reviews the use of large language models (LLMs) for social simulation and outlines the prevailing methodological pipeline, noting that current implicit natural-language agent specification approaches—while expressive and flexible—hinder rigor and comparability (Bail, 2024; Pang et al., 2025).
2.1 Believable Agents for Social Simulation
Believable Agents are interactive computational entities, originating as non-player characters in games (Bates, 1994). Early designs relied on rule-based systems or behavior trees to dictate their actions and decision-making processes (Loyall, 1997). Their observable behavior appears coherent, context-appropriate, and consistent with prior experiences, creating an “illusion of life” or human-like plausibility (Park et al., 2023; Anthis et al., 2025b; Banovic et al., 2023; Ehsan et al., 2024). The goal is not human-equivalent cognition but externally credible behavior: agents should perceive, plan, and act in ways that are explainable, temporally consistent, and aligned with social and physical norms (Goyal et al., 2024; Chandramouli et al., 2024; Liu et al., 2025a; Zhang et al., 2023).
Recent works use LLMs to create Believable Agents for social simulation (Bail, 2024; Wu et al., 2022; Anthis et al., 2025a; Hwang et al., 2025; Tjuatja et al., 2024; Prpa et al., 2024; Aubin Le Qu’er’e et al., 2024). For example, prior work (Park et al., 2023; Zhou et al., 2024) utilizes LLMs’ language understanding to prompt models with natural language persona descriptions, creating LLM-based agents that interact within simulated environments. These agents facilitate the study of emergent phenomena—such as cooperation, norm formation, and conflict—and heuristic behaviors like conformity (Liu et al., 2025b; Zhang et al., 2024; Shao et al., 2023a; Wu et al., 2025; Zhu et al., 2025). While these LLM-based Believable Agents broaden feasible questions and accelerate iteration by reducing labor, and mitigating ethical concerns (Bail, 2024; Hwang et al., 2025), they struggle with consistent, controllable behavior (Yang et al., 2025; Han et al., 2025)—a critical need for rigorous social science simulations (Bail, 2024).
In this work, we adopt the concept of Believable Agents and address prior rigor limitations by introducing CoBRA, a toolkit for systematic agent specification that yields consistent, controllable behavior.
2.2 Reproducibility of LLM-Based Methods in Social Science
Reproducibility is a critical concern in the use of LLMs for social science research, as the probabilistic nature of LLMs combined with varying implementation methods often leads to inconsistent and non-reproducible outcomes (Bail, 2024; Zhao et al., 2021).
One major issue is that different LLMs exhibit divergent behaviors even with the same prompts (Liu et al., 2023). Variations in training datasets (Tayo et al., 2023), model architectures (Jiang et al., 2024), and fine-tuning processes (Touvron et al., 2023) exacerbate this inconsistency, while the opacity of proprietary models further complicates cross-model comparisons and limits generalizability (Gudibande et al., 2023). Second, inference-time configurations significantly affect reproducibility. Sampling parameters like temperature settings influence output variability, with models differing in how peaky or flat their probability distributions are, affecting the generation coherence and diversity (Peeperkorn et al., 2024). Similarly, reasoning modes (Wei et al., 2022) (direct response vs. multi-step reasoning) introduce additional variability during inference, as different reasoning strategies can lead to divergent conclusions from the same input. Finally, post-training methods (e.g., RLHF (Ouyang et al., 2022)) also introduce inconsistencies across models, further complicating reproducibility. Addressing these challenges requires standardized methodologies and explicit control mechanisms to ensure reliable results.
In this work, we address reproducibility challenges by complementing the implicit natural language Agent Specification with an explicit approach that leverages structured, validated knowledge from classic social experiments to define agent behaviors, reducing variability and ensuring consistent outcomes.
2.3 Classic Social Experiments
Classic social experiments (French et al., 1959; Asch, 1951; Tversky & Kahneman, 1981; Milgram, 1963; Kahneman & Tversky, 2013; Wason, 1960; Lord et al., 1979; Goldstein et al., 2008; Milgram et al., 1969) provide structured and validated knowledge: long-replicated classic social experimental paradigms operationalize social-psychological constructs into clear manipulations and measurement protocols (Likert, 1932; Vazire & Mehl, 2008; Lilienfeld et al., 2017); the resulting behavioral patterns have been extensively validated and distilled into reusable, codified behavioral patterns (Unknown, 2015; Baron, 2007; Cialdini, 1993; Thaler & Sunstein, 2008).
As a concrete focus in this paper, we ground agent specification in cognitive biases (Unknown, 2015; Baron, 2007; Sumita et al., 2025)—a representative class of codified behavioral patterns—defined as systematic, predictable deviations from normative judgment and decision-making that arise from heuristics (Tversky & Kahneman, 1974; Simon, 1990). Prominent examples include Confirmation Bias (favoring information that confirms pre-existing beliefs, validated through the Wason Selection task (Wason, 1960) and Biased Information study (Lord et al., 1979)), Bandwagon Effect (adopting behaviors due to social conformity, demonstrated in Asch’s Line experiment (Asch, 1951) and the Hotel Towel study (Goldstein et al., 2008)), and so on. These cognitive biases, repeatedly validated in classic experiments, provide a rigorous foundation for specifying and evaluating LLM-based agent behaviors.
3 Understanding Agent Specification
This section presents a pilot experiment to empirically examine the limitations of implicit natural language specification. The results showed that current methods failed to produce consistent, cognitive bias-aligned behavior.
3.1 Method
Our key idea is to use a classical social experiment, the Asian Disease study (Tversky & Kahneman, 1981), as the study instrument. We created agents using different specification methods, powered these agents using different foundation models, and then analyzed their responses in the simulated social experiment.
Agent specifications. We evaluated two representative examples of implicit natural language agent specifications: a persona-based specification from (Park et al., 2023), which constructs agents with human-like profiles by incorporating demographics, personality traits, interests, and social relationships; and a role-based specification from (Chen et al., 2024b), which defines agents through task-oriented identities, assigning roles, domain expertise, and interaction protocols to guide task-specific behavior.
For the persona-based example (Park et al., 2023), we introduced three conditions: Common (an ordinary person), Economist (a persona with a professional background expected to make them less susceptible to framing effects and, therefore, less likely to prefer the positive frame compared to the Common persona), and Blank (no specification). We adopted the John Lin prompt from (Park et al., 2023) to represent the Common persona (see Appendix B for full prompt):
John Lin is a pharmacy shopkeeper at the Willow Market who loves to help
people. He is always looking for ways to make the process of getting
medication easier for his customers. John Lin is living with his wife,
Mei Lin, who is a college professor, and son, Eddy Lin, who...

### v3

# CoBRA: Programming Cognitive Bias in Social Agents Using Classic Social Science Experiments

@@ Abstract.
This paper introduces CoBRA [^1] This is the author’s accepted manuscript of a paper conditionally accepted to ACM CHI 2026. The final Version of Record is available via the ACM Digital Library., a novel toolkit for systematically specifying agent behavior in LLM-based social simulation. We found that conventional approaches that specify agent behavior through implicit natural-language descriptions often do not yield consistent behavior across models, and the resulting behavior does not capture the nuances of the descriptions. In contrast, CoBRA introduces a model-agnostic way to control agent behavior that lets researchers explicitly specify desired nuances and obtain consistent behavior across models. At the heart of CoBRA is a novel closed-loop system primitive with two components: (1) Cognitive Bias Index that measures the demonstrated cognitive bias of a social agent, by quantifying the agent’s reactions in a set of validated classic social science experiments; (2) Behavioral Regulation Engine that aligns the agent’s behavior to exhibit controlled cognitive bias. Through CoBRA, we show how to operationalize validated social-science knowledge (i.e., classical experiments) as reusable “gym” environments for AI—an approach that may generalize to richer social and affective simulations beyond bias alone.
AI for social science; HCI toolkit; social simulation; scientific reproducibility; large language models; representation engineering
$c$s: Computing methodologies Machine learningccs: Computing methodologies Natural language processingccs: Human-centered computing Collaborative and social computing
Refer to caption A three-part illustrative figure motivating the paper. The first part depicts the concept of framing effects in decision-making scenarios. The second part shows that agents specified only through implicit natural language descriptions exhibit inconsistent and unpredictable behaviors across different models, failing to reliably reflect the expected role-based differences. The third part introduces the proposed toolkit, which enables explicit and quantitative control of cognitive biases in social agents to produce more consistent and interpretable behaviors.
@@
Figure 1. Existing social simulation experiments often use implicit natural language descriptions to specify agent behaviors, which can lead to inconsistent and unpredictable outcomes. For example, Ⓐ a social simulation might expect agents playing economist roles to be less susceptible to framing effects than agents play the general population roles (Thaler and Sunstein, 2009); however, Ⓑ agents based on implicit natural language specifications often produce inconsistent behaviors across models, and the expected differences in behavior across roles (e.g., economists being less prone than laypeople) are not reliably observed. To address these challenges, Ⓒ, we introduce CoBRA, a novel toolkit that enables researchers to explicitly and quantitatively control the amount of cognitive biases exhibited in social agents’ observable behaviors.
1. Introduction
@@
A step-by-step diagram illustrating the closed-loop workflow of CoBRA, in which a researcher specifies a target framing-effect level for an agent, the system measures the agent’s bias using classic social science experiments such as the Asian Disease paradigm, and a behavioral regulation engine iteratively adjusts the agent through prompt engineering, activation-space control, or fine-tuning until the measured bias matches the target level.
Figure 2. Example closed-loop workflow of CoBRA. A social scientist aims to create an agent with a moderate framing effect (e.g., 2.6 on a 0–4 scale). ① She specifies the desired bias level in CoBRA alongside the natural language agent description. ② CoBRA measures the agent’s framing effect using validated classic social science experiments (e.g., the Asian Disease study (Tversky and Kahneman, 1981)). ③ If the measured bias deviates from the specification, the Behavioral Regulation Engine iteratively adjusts the agent—through prompt engineering, activation modifications, or fine-tuning—until the agent reliably demonstrates the target bias.
@@
The simulation of complex human behaviors through LLM-based agents is rapidly emerging as a powerful tool for social scientists (Bail, 2024; Park et al., 2023; Rahwan et al., 2019), offering new ways to test social theories (Anthis et al., 2025a; Wu et al., 2025) and explore sensitive or hard-to-study behaviors without real-world consequences (Hwang et al., 2025; Pang et al., 2025). A critical component of these simulations is specifying agent behaviors (Yang et al., 2025; Chandramouli et al., 2024; Goyal et al., 2024) to control how the agent responds to prompts. For example, Park et al. specified the attributes of one of the agents in their paper using the following natural language description: ”John Lin is a pharmacy shopkeeper at … who loves to help people. He is always looking for ways to make the process of getting medication easier for his customers…” (Park et al., 2023).
@@
These implicit natural language descriptions have become the dominant approach for specifying social agent behaviors and are used in nearly all social simulation experiments (Rath et al., 2025; Park et al., 2023; Zhou et al., 2024; Li et al., 2023a; Zheng et al., 2023; Chen et al., 2024a; Jia et al., 2024). In the descriptions, researchers often specify demographic descriptors (e.g., age, occupation, family structure), personality traits (e.g., Big Five (Gosling et al., 2003) or MBTI (Furnham, 1996)), personal interests (e.g., hobbies, preferences), social relationships (e.g., friends, family, colleagues), and professional or task-oriented identities (e.g., judge, programmer). These practices rest on a growing understanding that LLMs, having been trained on vast text corpora, possess both perceptual skills (Wang et al., 2024a; Liu et al., 2024) that enable them to detect subtle cues in implicit language (Gallegos et al., 2024), and strong role-playing capabilities (Park et al., 2023)—drawing on patterns that link behaviors to human characteristics (Fast et al., 2016; Shao et al., 2023).
@@
However, the research community has little empirical understanding of whether this approach is sufficient for reliable social simulations. For example, would the conclusion of the social simulation change if we change the foundation models, the prompts, or the temperatures? As a consequence, the conclusions of most LLM-based social simulations often are broad qualitative impressions—such as noting that a simulation is plausible and exhibits an echo chamber similar to the real world (Wang et al., 2024b; Zheng and Tang, 2024; Moore et al., 2024).
@@
In this paper, we first conducted a pilot experiment to empirically examine the limitations of implicit natural language specification (Section 3). We designed four agent specifications by adapting examples from prior work (Chen et al., 2024a; Park et al., 2023), implemented them across four foundation models, and evaluated their responses using a classic social science experiment, the Asian Disease study (Tversky and Kahneman, 1981). The results show that (1) the same specification produced inconsistent behaviors across models, and (2) implicit specification did not reliably produce the expected behaviors (e.g., trained experts like economists are less prone to the framing effect (List, 2003)).
@@
We then developed CoBRA [^2] The source code is publicly available at https://github.com/AISmithLab/CoBRA—the Cognitive Bias Regulator for Social Agents (Figure 1)—a novel toolkit that enables researchers to explicitly and quantitatively control the amount of cognitive biases exhibited in social agents’ observable behaviors. Compared to implicit natural language specifications, CoBRA offers two key advantages. First, it enhances reproducibility of agent behaviors across models, which is critical given that access to specific foundation models may change over time (Chen et al., 2024b). Second, it provides explicit and quantitative control over agent behaviors, rather than relying on unpredictable model behaviors (Sclar et al., 2023). For example, a scientist can directly specify that an agent should be less susceptible to framing effects.
@@
At the heart of CoBRA is a novel closed-loop system primitive that continuously measures the amount of cognitive biases demonstrated in agents’ responses to a set of validated classic social science experiments and adjusts the agent specifications accordingly. Classic experiments provide validated and structured protocols: they define precise stimuli, expected behavioral distributions, and well-documented bias effects. For instance, the Asian Disease study (Tversky and Kahneman, 1981) has been extensively replicated and is widely regarded as a canonical demonstration of the framing effect: when identical outcomes are framed in terms of lives saved versus lives lost, human participants systematically shift between risk-averse and risk-seeking choices. Within CoBRA, such experiments function as calibration tasks: the system first measures whether an agent demonstrates the expected bias by observing its responses in these experiments and then adjusts the agent’s specification to align it with the target behavioral profile (Figure 2).
@@
Our implementation of CoBRA has two components: (1) the Cognitive Bias Index that measures the cognitive bias of a social agent by quantifying the agent’s reactions in a set of validated classic social science experiments; (2) the Behavioral Regulation Engine that aligns the agent’s behavior to demonstrate controlled cognitive bias. The Cognitive Bias Index contains eight classic social experiments mapped to four representative biases—Authority Effect, Bandwagon Effect, Confirmation Bias, and Framing Effect. The Behavioral Regulation Engine employs three approaches (prompt engineering (Rath et al., 2025; Zhuo et al., 2024), representation engineering (Zou et al., 2023; Rimsky et al., 2024; Todd et al., 2024), and fine-tuning (Ouyang et al., 2022; Rafailov et al., 2023; Hu et al., 2022; Ilharco et al., 2023)) to align agent behaviors.
@@
We position CoBRA as an HCI toolkit (Nebeling, 2017) that empowers social scientists to create new types of social simulations. We evaluated CoBRA using two common strategies for assessing HCI toolkits (Ledo et al., 2018): Technical Benchmark and Demonstration. In the benchmark, we show that (1) agents specified through CoBRA exhibit consistent behavioral tendencies regardless of the foundation model, sampling temperature, or reasoning mode (Section 7.1); (2) CoBRA enables precise and predictable control under both open-weight and API-only settings (Section 7.2). In the Demonstration, we show how social scientists can use CoBRA to simulate the Emotional Contagion experiment (Kramer et al., 2014), where the participant’s social influence susceptibility (i.e., bandwagon effect bias) can be manipulated (Section 8). The simulated experiment validates a classic social science theory: agents exposed to varying levels of negative posts exhibited predictable sentiment shifts based on their bias levels programmed by CoBRA, with higher Bandwagon Effect levels leading to stronger emotional contagion.
@@
This work makes two main contributions. First, we introduce CoBRA, a novel HCI toolkit that enables researchers to explicitly and quantitatively control the amount of cognitive biases exhibited in social agents’ observable behaviors, improving reproducibility of LLM-based social simulations and allowing explicit, quantitative alignment. Second, we contribute a closed-loop system primitive that grounds agent behaviors in validated social science experiments, using them as primitives to measure and regulate cognitive biases, and that can in principle extend to other social and affective phenomena.
Scope and limitations. We adopt a functional definition of cognitive bias: a construct operationalized in validated social-science experiments and identified through measurable patterns in agents’ observable responses (Tversky and Kahneman, 1974). Aligning an agent’s behavior with these patterns does not imply that the agent’s internal representations instantiate the corresponding human bias.
2. Related Work
CoBRA is built on a simple idea: use classic social science experiments as calibration tasks to control agent behavior. This design can improve agent behavior reproducibility across models and allow users to actively control agent behaviors. We situate CoBRA within three strands of related work: (1) LLM-based agent-based modeling, (2) aligning model and agent behavior, and (3) unpredictable user interfaces in LLMs.
2.1. LLM-based Agent-based Modeling
Agent-based modeling (ABM) simulates the actions and interactions of many autonomous agents to explore “what-if” scenarios that would be difficult, expensive, or unethical to test in the real world (Railsback and Grimm, 2019; Bonabeau, 2002). ABMs have proven useful across various domains, including public health (Tracy et al., 2018), economics (Axtell and Farmer, 2025), and disaster response (Hawe et al., 2012). For example, Schelling’s spatial proximity model of segregation (Pancs and Vriend, 2007) treats households as agents on a grid who prefer to live near a minimum fraction of similar neighbors; even mild individual preferences, when simulated at scale, produce starkly segregated neighborhoods, illustrating how simple local rules can generate undesirable macro-level patterns. However, a major limitation of conventional ABM is the limited expressiveness of its agents. Most models rely on overly simplistic, rule-based behaviors that cannot capture the nuanced, adaptive decision-making of real-world individuals (Bonabeau, 2002; Bates, 1994).
Recent studies suggest that LLMs may help overcome this limitation, as LLMs possess both perceptual skills (Bail, 2024; Tjuatja et al., 2024; Aubin Le Qu’er’e et al., 2024; Wu et al., 2025; Hwang et al., 2025) and role-playing capabilities (Park et al., 2023; Anthis et al., 2025a; Prpa et al., 2024; Liu et al., 2025b; Zhang et al., 2024; Shao et al., 2023; Zhu et al., 2025; Zhou et al., 2024). For example, Park et al. built a small-town “simulacrum” where dozens of LLM-based agents interacted over time (Park et al., 2023). As the simulation unfolded, agents developed daily routines and even exhibited emergent group behaviors, such as organizing and discussing attendance at a party (Park et al., 2023). However, there is ongoing debate about whether LLMs can reliably simulate human behaviors (Bail, 2024; Zhao et al., 2021; Yang et al., 2025; Han et al., 2025; Koo et al., 2023; Anthis et al., 2025b).
CoBRA bridges these two lines of work by introducing a new form of agent specification that uses cognitive biases (Cheung et al., 2025; Echterhoff et al., 2024; Cosmides and Tooby, 1994; Boonprakong et al., 2025; Haselton et al., 2015; Sumita et al., 2025) as the primary control abstraction. Like conventional agent-based models, it offers a simple interface for specifying agent behavior. At the same time, it leverages modern LLMs to implement agents more expressively.
2.2. Model/Agent Behavior Aligning
Prior work on alignment has focused on steering model behavior to better reflect human preferences and normative constraints (Ouyang et al., 2022; Shi et al., 2024; Rimsky et al., 2024; Goyal et al., 2024; Choenni et al., 2024; Lyu et al., 2024). More recently, researchers have begun to apply similar ideas at the agent level. These approaches typically focus on learning a reward model that scores outputs based on how well they match the desired output. For example, in human-feedback alignment, annotators rank multiple candidate responses, and the model is then optimized—via reinforcement learning or direct preference optimization—to prefer higher-ranked responses (Ouyang et al., 2022; Rafailov et al., 2023; Li et al., 2025). Recent agent-based systems further constrain behavior by encoding high-level “constitutional” rules (e.g., be cooperative, follow safety guidelines) and optimizing agents to follow these rules during interaction (Chen et al., 2025; Arditi et al., 2024; Zhuo et al., 2024). Cognitive biases are operationalized as systematic patterns in observable judgments and choices (Tversky and Kahneman, 1974; Liu et al., 2025b). Classic work on heuristics and decision making defines bias through consistent deviations in decision behavior across controlled tasks such as probabilistic reasoning and framing (Milgram, 1963; Asch, 1951; Tversky and Kahneman, 1981; Nisbett and Wilson, 1977; Lord et al., 1979; Wason, 1960). These experimentally validated paradigms provide a principled basis for treating cognitive bias as a behavioral property that can be measured and systematically studied.
Building on this view, the core idea behind CoBRA is to harness classic social science experiments to generate synthetic training (or prompt calibration) data. Decades of work have replicated and refined these experiments, publishing detailed protocols that link experimental manipulations, observed behaviors, and underlying cognitive biases. We translate experimental conditions into prompts and canonical behavioral responses into targets, then align the model on this synthetic dataset to induce controlled levels of bias in agent behavior, consistent with the approximately linear structure of LLM representations discussed in recent work (Park et al., 2024). This parallels recent Embodied AI work that uses physics simulators to generate synthetic data (Szot et al., 2021), but here the “simulator” is grounded in empirical social experiments rather than physical laws.
2.3. Unpredictable User Interfaces in LLMs
CoBRA also draws on HCI theories of mental models (Norman, 2014) and gulfs between user intentions and system behavior (Norman, 1986). Prior work has highlighted that unpredictable black-box systems make for poor user interfaces because users cannot reliably anticipate how small changes in input will affect output (Yang et al., 2025; Banovic et al., 2023; Ehsan et al., 2024; Subramonyam et al., 2024). LLM-based tools often exhibit this problem (Agrawala, 2023; Subramonyam et al., 2024). For example, changing a bias from 0.1 to 0.2 could change agent behavior far more than a shift from 0.2 to 0.5 (see Figure 10). Switching the model versions or types may also shift behaviors (Chen et al., 2024b; Liu et al., 2023; Santurkar et al., 2023; Jiang et al., 2024; Touvron et al., 2023; Gudibande et al., 2023; Peeperkorn et al., 2024; Wei et al., 2022). Recent research has begun to address this by adding interactivity (Brade et al., 2023; Zhang et al., 2023; Wu et al., 2022; Goel et al., 2024; Liu et al., 2025a), for example, through interfaces that let users iteratively refine prompts, inspect alternative responses, or visualize how prompt edits affect model behavior.
CoBRA explores a complementary direction by introducing a calibrated control proxy that better matches users’ mental models of conventional interfaces. Rather than manipulating prompts or model settings directly, which often yields unpredictable behavior, users specify an interpretable cognitive bias index, which CoBRA translates into underlying alignment parameters. This translation layer is calibrated so that the mapping is monotonic, smooth, and expressive: small adjustments to the index (e.g., increasing a bias level) reliably produce small, predictable shifts in agent behavior, much like turning a dial or moving a slider in a traditional interface.
3. Understanding Agent Specification
We first conducted a pilot experiment to assess the limitations of implicit natural language specification.
3.1. Method
We used Asian Disease (Tversky and Kahneman, 1981) as the study instrument. Asian Disease is a classic social experiment for testing the framing effect. There are two validated insights associated with this experiment: (1) Participants tended to make more risk-averse decisions when outcomes were positively framed (e.g., “200 people will be saved”), but exhibited greater risk-seeking behavior when the same outcomes were presented as losses (e.g., “400 people will die”) (Tversky and Kahneman, 1981); (2) Domain expertise (e.g., expertise in economics) can attenuate susceptibility to framing effects (Thaler and Sunstein, 2009), which we treat as a tendency observed in prior work rather than a universal property of all economists or all contexts, and use as a hypothesis about how an “Economist” persona might behave under framing. We developed 15 scenario prompts by adapting published experiment descriptions (Tversky and Kahneman, 1981; Kahneman and Tversky, 2013), each with varying positive and negative framings. For example,
You are presented with a scenario where a disease outbreak is
expected to kill 600 people unless a program is implemented.
Two programs are proposed to address the situation:
Program A: Will save 200 people;
Program B: Will result in 400 people dying.
We then asked the agent to answer a five-point Likert-style multiple-choice question, with options ranging from a strong preference for the positively framed, through intermediate and neutral choices, to a strong preference for the negatively framed.